### v3

# Translating Images to Road Network: A Sequence-to-Sequence Perspective

Abstract
The extraction of road network is essential for the generation of high-definition maps since it enables the precise localization of road landmarks and their interconnections. However, generating road network poses a significant challenge due to the conflicting underlying combination of Euclidean (e.g., road landmarks location) and non-Euclidean (e.g., road topological connectivity) structures. Existing methods struggle to merge the two types of data domains effectively, but few of them address it properly. Instead, our work establishes a unified representation of both types of data domain by projecting both Euclidean and non-Euclidean data into an integer series called RoadNet Sequence. Further than modeling an auto-regressive sequence-to-sequence Transformer model to understand RoadNet Sequence, we decouple the dependency of RoadNet Sequence into a mixture of auto-regressive and non-autoregressive dependency. Building on this, our proposed non-autoregressive sequence-to-sequence approach leverages non-autoregressive dependencies while fixing the gap towards auto-regressive dependencies, resulting in success in both efficiency and accuracy. We further identify two main bottlenecks in the current RoadNetTransformer on a non-overfitting split of the dataset: poor landmark detection limited by the BEV Encoder and error propagation to topology reasoning. Therefore, we propose Topology-Inherited Training to inherit better topology knowledge into RoadNetTransformer. Additionally, we collect SD-Maps from open-source map datasets and use this prior information to significantly improve landmark detection and reachability. Extensive experiments on the nuScenes dataset demonstrate the superiority of RoadNet Sequence representation and the non-autoregressive approach compared to existing state-of-the-art alternatives. Our code is publicly available at open-source https://github.com/fudan-zvg/RoadNetworkTRansformer.
Index Terms:
Road Network, Transformer, Sequence-to-sequence.
I Introduction
Figure 1: High-definition Road Network Topology contains Euclidean data: locations of landmarks and shapes of curves and non-Euclidean data: road topology. A special sequence, RoadNet Sequence, is proposed as a unified representation of both domains. Then we use a Non-Autoregressive sequence-to-sequence approach to extract RoadNet Sequence from multi-camera input efficiently and accurately.
With the rising prevalence of self-driving cars, a deep knowledge of the road structure is indispensable for autonomous vehicle navigation [1, 2, 3, 4]. Road network [5, 6, 7] extraction is required to estimate highly accurate road landmark locations, centerline curve shapes, and road topological connection for self-driving vehicles. However, the ability to understand road network in real-time using onboard sensors is highly challenging.
In the literature, road network is differentiated from methods that focus solely on grid-like Euclidean data [8, 9] such as lane detection [10, 11, 12] or BEV semantic understanding [13, 14, 15, 16]. Instead, road network emphasizes a more comprehensive understanding of both Euclidean domain and non-Euclidean domain. As shown in Figure 1, accurate road landmark locations such as crossroads, stop-lines, fork-roads, and the shape of centerline curves pertain to the Euclidean domain, whereas road topology belongs to the non-Euclidean domain. In mathematical terms [9, 8], Euclidean data is defined in $\mathbb{R}^{2}$. On the other hand, non-Euclidean data, such as graphs that indicate connectivity among nodes, will lose crucial information if projected to $\mathbb{R}^{n}$, such as edge curve information.
Unfortunately, existing attempts [17, 18] to extract road network using onboard sensors have not been able to achieve a harmonious integration between Euclidean and non-Euclidean data. STSU [17] divides the road network construction into two stages: center-lines detection and center-line connectivity reasoning—which ignores the cooperation between Euclidean and non-Euclidean domains. TPLR [18] uses a Transformer or Polygon-RNN [19] to combine topology reasoning with center-line localization, but the embedded conflicts between Euclidean data representations and connectivity representations undermine the model performance.
In this work, we propose that the dilemma in existing works arises due to the absence of a unified representation of both Euclidean and non-Euclidean data. Instead, we introduce a Euclidean-nonEuclidean unified representation with merits of losslessness, efficiency and interaction. The unified representation, named as RoadNet Sequence, projects both Euclidean and non-Euclidean aspects of road network to integer series domain $\mathbb{Z}^{n}$. The “losslessness” aspect is ensured by establishing a bijection from road network to RoadNet Sequence. “Efficiency” is achieved by limiting RoadNet Sequence length to the shortest $\mathcal{O}(|\mathcal{E}|)$ (where $\mathcal{E}$ is the set of all centerlines), through a specially designed topological sorting rule. “Interaction” reveals the interdependence between Euclidean and non-Euclidean information within a single sequence.
Based on the auto-regressive dependency of topological sorting, we leverage the sequence-to-sequence generation power of Transformer [20, 21, 22] to understand RoadNet Sequence from onboard round-view cameras, called RoadNetworkTRansformer (RNTR).
However, the varying order of node and edge generation results in different auto-regressive dependencies in sequence-to-sequence generation. In practice, we examine two types of dependencies: coupled RoadNet Sequence and decoupled RoadNet Sequence. The coupled RoadNet Sequence generates each node and immediately outputs the edge connected to it, whereas the decoupled RoadNet Sequence first generates all nodes and then outputs the edges connected to each node.
Moreover, the decoupled RoadNet Sequence can be upgrated to accelerate the slow auto-regressive process. Our observation is that the dependency of road network can also be decoupled into a semi-autoregressive format, that retains auto-regressive functionality within local contexts while simultaneously conducting multiple generations in parallel. This semi-autoregressive model is called Semi-Autoregressive RoadNetTransformer (SAR-RNTR). This approach not only accelerates the inference speed by 6 times, but it also significantly boosts the accuracy to a new level based on the better dependency modeling. Going beyond SAR-RNTR, we employ a masked training technique on SAR-RNTR to mimic the remaining auto-regressive dependency through iterative prediction [23]. This gives rise to our Non-Autoregressive RoadNetTransformer (NAR-RNTR) model, which achieves real-time inference speed ($47\times$ faster) while maintaining the high performance.
To evaluate road network extraction quality, apart from inheriting the former metrics [17] based on lane detection, we instead design a family of metrics directly based on the definition of the road network – Landmark Precision-Recall and Reachability-Precision-Recall – to evaluate (i) road landmarks localization accuracy, and (ii) path accuracy between any reachable landmarks.
We make the following contributions: (i) we introduce RoadNet Sequence, a lossless, efficient and unified representation of both Euclidean and non-Euclidean information from road network. (ii) We propose a Transformer-based RoadNetTransformer which can decode RoadNet Sequence from multiple onboard cameras. (iii) By decoupling auto-regressive dependency of RoadNet Sequence, our proposed Non-autoregressive RoadNetTransformer accelerates the inference speed to real-time while boosting the accuracy with a significant step from the auto-regressive model. (iv) Extensive experiments on nuScenes [24] dataset validate the superiority of RoadNet Sequence representation and RoadNetTransformer over the alternative methods with a considerable margin.
In the extension, we address the major bottlenecks of sequence-to-sequence models for road network extraction. During our exploration, we split the nuScenes [24] dataset into non-overlapping scenes to evaluate the non-overfitting performance. Unlike before [25], we also ensure that different daytime and weather conditions are equally distributed in the training and evaluation sets. The bottlenecks arise from poor road perception, and the errors subsequently propagate to the understanding of the topology. To mitigate error propagation in topology learning, we propose a new training strategy called Topology-Inherited Training (TIT), which inherits topology knowledge from a large dataset trained model. TIT first trains a high-performance SDMap-RoadNetTransformer (SD-RNTR) by using bird’s-eye-view (BEV) maps as input instead of round-view images to simplify Euclidean data perception. Additionally, we use the sweep data in nuScenes [24], which is nine times larger than the traditional nuScenes dataset, to learn better non-Euclidean perception. We then distill the BEV representation from the encoder pretrained with lidar maps to the traditional round-view image input BEV Encoder. Finally, we assemble the BEV Encoder with round-view image input and RoadNetTransformer, boosting performance across different modality inputs. Apart from enhancing non-Euclidean data learning, the distillation of round-view image BEV representation learning can also improve the performance of Euclidean data.
Figure 2: Comparison between Navigation Map (e.g., Google Maps [26] or OpenStreetMap [27]), SD-Map (extracted from OpenStreetMap), and Lidar Map (collected by the nuScenes [24] dataset).
Although the TIT partly solve the bottlenecks, the further use of standard definition maps (SD-Maps) from daily navigation applications like Google Maps provides essential prior information about basic landmarks and topology. As shown in Figure 1, unlike road network, which contains lane-level and centimeter-level accuracy information, SD-Maps offer only road-level information, indicating road connections with meter-level accuracy. Consequently, SD-Maps cannot provide accurate lane numbers or precise landmark locations. We collect the corresponding SD-Map data from OpenStreetMap [27] (similar to Google Map) and manually align it with the nuScenes [24] dataset. By treating SD-Maps as a prompt learning problem and integrating them into the Transformer model, we achieve significant performance improvements without altering the model structure or requiring complex data processing.
In conclusion, we make the following extension: (i) We examine the advantages of decoupled RoadNet Sequence; (ii) We propose a brand-new training strategy, Topology-Inherited Training, to mitigate error propagation in topology learning, enhancing both Euclidean and non-Euclidean data understanding; (iii) We introduce SDMap-RoadNetTransformer, which realizes the inheritance of topology knowledge across modalities; (iv) We collect SD-Map data from OpenStreetMap [27] and treat SD-Maps as a prompt learning problem without altering the model structure or complex data processing.
II Related work
Vision-based ego-car (BEV) feature learning. A rising tendency is conducting self-driving downstream tasks [28, 15, 16, 29, 13, 14, 30, 31, 32, 33] under ego-car coordinate frame. To learn ego-car feature learning from onboard cameras,  [34, 35] project image features to ego-car coordinate based on depth estimation. OFT [28], LSS [13] and FIERY [30] predict depth distribution to generate the intermediate 3D representations. [14, 25, 15, 16] resort neural network like Transformer to learn ego-car feature without depth. In order to learn the topology of the road network under the ego-car coordinate frame, we use the straightforward method of applying LSS [13, 29] to extract ego-car features from the multiple onboard cameras.
Road network extraction. Researchers have explored the utilization of DNNs to decode and recover maps from aerial images and GPS trajectories [36, 37]. Moreover, STSU [17] first detects centerline from front-view image with Transformer and then predicts the association between centerline with MLP layer, followed by a final merge to estimate road network. Based on STSU, TPLR [18] introduces minimal cycle to eliminate ambiguity in its connectivity representation. Existing methods spend great effort to deal with the problem in the non-Euclidean domain but ignore the cooperation between Euclidean and non-Euclidean. Previous approaches, such as STSU [17] and TPLR [18], focus on solving the problem within the non-Euclidean domain, typically handling road network connectivity through centerline detection or minimal cycles. While these methods tackle topological reasoning and road connectivity, they often neglect the interdependencies between Euclidean and non-Euclidean information. For example, STSU detects centerlines using Transformer-based models and then predicts the association between centerlines through an MLP, but it does not effectively merge road geometry with topological structure. TPLR addresses connectivity ambiguities by introducing minimal cycles, but it does not fully integrate Euclidean data with non-Euclidean topological data. In contrast, our work proposes a unified representation that integrates both Euclidean and non-Euclidean road network information into a single framework, overcoming the limitations of previous methods. Instead of treating these domains separately, as done by STSU [17] and TPLR [18], we introduce a new method that projects both Euclidean and non-Euclidean data into an integer series called RoadNet Sequence. By doing so, we create a unified, lossless, and interactive representation that allows for more accurate modeling of road networks. Further, we introduce a sequence-to-sequence Transformer model that decouples the dependency of RoadNet Sequence into a mixture of auto-regressive and non-autoregressive dependencies. This innovative approach allows our model to leverage the benefits of both dependencies, ensuring high efficiency and accuracy in predicting road network properties. Our proposed non-autoregressive approach significantly improves the efficiency of training and inference while maintaining the model’s predictive accuracy.
Non-Autoregressvie generation. Non-Autoregressive generation for Neural Machine Translation (NMT) [38] has been proposed as a solution to speed up the one-by-one sequential generation process of autoregressive models. [39, 40, 41] utilized knowledge distillation to help the NAT model capture target sequence dependency. [23, 42, 43, 44] proposed iterative models that refine the output from the previous iteration or a noised target in a step-wise manner. [45] kept the auto-regressive approach for global modeling, but introduced the parallel output of a few successive words at each time step. [46, 47, 48, 49] used latent variables as intermediates to reduce the dependency on the target sequence. The current state of non-autoregressive NMT models is far from perfect, with a significant gap in performance compared to their auto-regressive counterparts [50]. But in the case of our RoadNet Sequence, the auto-regressive dependency can be decoupled, leading to both acceleration and improved performance.
III Method
Figure 3: Top illustrates the transformation process from a Directed Acyclic Graph to a Directed Forest for RoadNet Sequence. The red boxes enclose all merge-points that have non-unique parents. We replicate theirs parents (except the first parent) to their children (cyan nodes) and delete the corresponding edges. Middle demonstrates a bijection between the edges and vertices ($\dagger$}: excluding the Ancestors). Bottom lists the topological sorting result of vertices and presents six integers for each vertex.
Figure 4: Left shows two major components of Semi-Autoregressive RoadNetTransformer: Key-point Transformer decoder to detect key-points and Parallel-seq Transformer Decoder to generate Semi-Autoregressive RoadNet Sequence. “SAR RoadNet Seq” stands for Semi-Autoregressive RoadNet Sequence. Right decouples self-attention of Parallel-seq Transformer Decoder layer into Intra-seq and inter-seq self-attention.
In this section, we will introduce: (i) mathematics modeling of road network, (ii) RoadNet Sequence, (iii) architecture of RoadNetTransformer, (iv) metrics.
III-A Mathematics Modeling of Road Network
The road network comprises road landmarks (e.g., crossroad, stop-lines, fork-point, merge-point) and centerlines connecting among them [17, 18]. Given that traffic on each lane is always one-way, the graph can be formulated as a Directed Acylic Graph (DAG), i.e., $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where vertex set $\mathcal{V}$ is the set of all road landmarks and edge set $\mathcal{E}$ is the set of all centerlines. Each vertex $v=(v_{x},v_{y},v_{c})\in\mathcal{V}$ contains two properties: (i) location, i.e., $(v_{x}\in\mathbb{R},v_{y}\in\mathbb{R})$, (ii) category, i.e., $v_{c}$. As shown in Figure 3, we divide vertices into 4 categories which will be introduced later. Each edge $e=(e_{s},e_{t},e_{px},e_{py})\in\mathcal{E}$ contains two properties: (i) source $e_{s}\in\mathcal{V}$ and target $e_{t}\in\mathcal{V}$ vertex of the edge, (ii) Bezier middle control point $(e_{px}\in\mathbb{R},e_{py}\in\mathbb{R})$ [17] controlling the shape of edge curve. By and large, the Euclidean data of $\mathcal{G}$ contains $v_{x},v_{y},e_{px},e_{py}$ while the non-Euclidean data contains $v_{c},e_{s},e_{t}$.
Figure 5: The sequence of a decoupled RoadNet Sequence comprises two components: the vertex sequence and the edge sequence. The vertex sequence consists of vertex coordinates arranged in depth-first search order and is encoded by their coordinates. For each vertex, the edges starting from it are listed in edge encoding in order. Finally, the vertex sequence and edge sequence are concatenated with special tokens to indicate the start and end.
III-B Coupled RoadNet Sequence
In this section, we present RoadNet Sequence to integrate Euclidean and Non-Euclidean data into a unified representation. The projection from road network to coupled RoadNet Sequence includes: (i) from DAG to Directed Forest, (ii) topological sorting of Directed Forest and (iii) sequence construction.
[…]
As shown in the top of Figure 3, the transformation happens for all the merge-points on the road, i.e., $\{v\in\mathcal{V}\ |\ id(v)>1\}$ whose parents are non-unique. We replicate their parents (except the first parent) to their children and delete the corresponding edges. Then we assign these created vertices with a specific label Clone so that the direction of the edge between the merge-points and themselves will be reverted during recovery.
Topological sorting of Directed Forest. We use Depth-first search (DFS) to obtain the sequential order of vertices $\mathcal{V}_{f}$ in the Directed Forest, i.e., $v_{1},v_{2},\cdots,v_{n}]$. But the Clone point labeled in the last paragraph will not be traversed. Instead, as shown in the bottom of Figure 3, Clone point will be traversed following its original point.
Thanks to the bijection $f$ between $\mathcal{E}_{f}$ and $\{v\in\mathcal{V}_{f}\ |\ id(v)=1\}$, we can build pairs for all edges and vertices, i.e., $(v,f(v))$ for $v\in\{v\in\mathcal{V}_{f}\ |\ id(v)=1\}$ and $(v,\text{None})$ for $v\in\{v\in\mathcal{V}_{f}\ |\ id(v)=0\}$. These pairs construct the road network with no redundancy.
Non-unique sorting. In the application of Depth-First Search (DFS), the sequence in which points under the same parent are searched first can result in non-unique sorting. We investigate two strategies to address this: a random ordering strategy and an ordering based on coordinates, specifically prioritizing points closest to the Bird’s Eye View (BEV) front right.
[…]
There exist 4 cases to determine the category $v_{c}$: (i) if $v_{i}$ is the root of the tree, the category $v_{c}$ is set as Ancestor, the index of its parent $v_{d}$ is set as None, its coefficient is ignored. (ii) if $v_{i}$ is the first child of its parent, the category $v_{c}$ is set as Lineal, its $v_{d}$ is ignored for its parent is exact $v_{i-1}$. (iii) if $v_{i}$ is not the first child of its parent, the category $v_{c}$ is set as Offshoot, its $v_{d}$ is its parent’s index. (iv) if $v_{i}$ is the cloned child, the category is set as Clone, its $v_{d}$ is its original child. Integer representation of $v_{c}$ is Ancestor: 0, Lineal: 1, Offshoot: 2, Clone: 3.
Details of sequence construction. The discretization of $v_{x},v_{y}$ is simply truncating the integer part. Integer representation of $v_{c}$ is Ancestor: 0, Lineal: 1, Offshoot: 2, Clone: 3. Discretizing $e_{px}$ and $e_{py}$ can be challenging since the Bezier control points may exceed the Bird’s Eye View (BEV) range, and their values may become negative. As a solution, we discretize $e_{px}$ and $e_{py}$ by applying the int function to $(e_{px}+10)$ and $(e_{py}+10)$, respectively, to avoid negative values.
Analysis. The proposed RoadNet Sequence possess the merits of losslessness, efficiency and interaction. The losslessness of RoadNet Sequence is guaranteed by establishing a bijection between edges and vertices (excluding the root) of trees. Our RoadNet Sequence has a complexity of $\mathcal{O}(|\mathcal{V}_{f}|)=\mathcal{O}(|\mathcal{E}|)$, making it the most efficient. A mixture of Euclidean and non-Euclidean data within a local 6-integer clause facilitates full interaction between the Euclidean and non-Euclidean domains.
The transformation from a Directed Acyclic Graph (DAG) to a Directed Forest shares similarities with the lookahead method in Transformers, especially in capturing sequential dependencies and enhancing future predictions. However, the key difference lies in the methodology: while the lookahead method in Transformers uses dynamic attention to model dependencies, our DAG-to-Directed Forest transformation simplifies the graph structure to reduce redundancy, making the prediction process more efficient. Additionally, while Transformers rely on embeddings and attention, our method uses a predefined structural transformation tailored for road network sequence prediction.
RoadNet Sequence also possess the auto-regressive dependency. Since Depth-First search of Trees is always topological sorting, vertices only depend on the previous generated vertices. Also, our 6 integers come in the order of vertices location, vertices category, index and curve coefficient also preserving the auto-regressive assumption.
Sequence embedding. Each vertex-edge pair is represented by 6 integers. To prevent embedding conflicts between the 6 integers, we divide them into separate ranges. As a default, we set the embedding size to $576$, which is sufficient to accommodate all the integer ranges.
III-C Decoupled RoadNet Sequence
Sequence construction. As shown in Figure 5, the sequence of a decoupled RoadNet Sequence comprises two components: the vertex sequence and the edge sequence.
The vertex sequence consists of vertices coordinates arranged in the topological sorting order as discussed before. Since lane graphs encompass varying vertex quantities, the resulting sequences will exhibit distinct lengths. In order to demarcate the conclusion of a vertex sequence, we introduce an <EOV> (End of Vertex Sequence) token.
An edge can be represented as a set of the source vertex, the target vertex, and the Bezier middle control point. For each vertex in DAG, We denote the target vertex using the sequential index of its child nodes within the vertex sequence. The source vertex of each edge is determined by arranging the edge in the order of its source vertex in the vertex sequence.
[…]
Besides, we use <Start> token to indicate the beginning of the whole sequence. The <N/A> token is utilized to separately pad the vertex sequence and edge sequence, which guarantees that both of them maintain a consistent length.
Analysis The proposed decoupled RoadNet Sequence also possess the merits of losslessness, efficiency and interaction. The losslessness of RoadNet Sequence is guaranteed by establishing a bijection between edges and vertices (excluding the root) of trees. The decoupled RoadNet Sequence has a complexity of $\mathcal{O}(|\mathcal{V}|)+\mathcal{O}(|\mathcal{E}|)=\mathcal{O}(|\mathcal{E}|)$, making it more efficient. A mixture of Euclidean and non-Euclidean data within the sentence also facilitates interaction between the Euclidean and non-Euclidean domains.
III-D Input and Target Sequence Construction
Sequence embedding. Each vertex-edge pair is represented by 6 integers. To prevent embedding conflicts between the 6 integers, we divide them into separate ranges which is shown in Table I. As a default, we set the embedding size to $576$, which is sufficient to accommodate all the integer ranges.
TABLE I: Embedding range of different integers.
Item Range
$v_{x},v_{y}$ $0\sim 199$
$v_{c}$ $200\sim 249$
$v_{d}$ $250\sim 349$
$e_{px},e_{py}$ $350\sim 569$
noise category $570$
EOS $571$
Start 57572
n/a $573$
Synthetic noise objects technique. The input sequence of RoadNet Sequence starts with a start token and the target sequence ends with an EOS token. The EOS token makes the model know where the sequence terminates, but the experiments have shown that it tends to cause the model to stop predicting early without getting the complete sequence. Inspired by [22], we use a similar sequence augmentation technique to alleviate the problem called the synthetic noise objects technique [22]. The technique composes of sequence augmentation and sequence noise padding. The sequence augmentation adds noise to the position of landmarks and the coefficient of centerlines in input sequence. Whereas, sequence noise padding is a padding technique. For input sequences, we generate synthetic noise vertices and append them at the end of the real vertices sequence. Each noise vertex includes random locations($v_{x}$, $v_{y}$), category($v_{c}$), index of parent($v_{d}$) and Bezier curve coefficient($e_{px}$, $e_{py}$). As for the target sequence, the EOS token is added to the end of the real vertices sequence. We set the target category($v_{c}$) of each noise vertex to a specific noise class(different from any of the ground-truth labels), and the remaining components($v_{x}$, $v_{y}$, $v_{d}$, $e_{px}$, $e_{py}$) of the noise vertex to the ”n/a” class, whose loss is not calculated in the back-propagation.
However, we only use sequence noise padding as sequence augmentation has been shown to cause a decrease in performance. The introduced modifications of the synthetic noise objects technique are illustrated in Figure 6.
The padding rules of Semi-Autoregressive RoadNet Sequence are much the same as auto-regressive RoadNet Sequence. As mentioned in the main submission, we pad the 2-dimensional Semi-Autoregressive RoadNet Sequence to $y_{1,1},y_{1,2},\cdots,y_{1,L}],\cdots,[y_{M,1},y_{M,2},\cdots,y_{M,L}]]$, where $L$ is the maximum length of each sub-sequence and $M$ is the number of sub-sequences. The valid sub-sequences begin with a key-point. For each valid sub-sequence, we follow the same padding rules of RoadNet Sequence, except there isn’t a start token in an input sub-sequence because of the Key-point Prompt. We set the other sub-sequences to the ”n/a” class making the loss of these sub-sequences without a key-point not calculated.
Figure 6: An illustration of synthetic noise objects technique [22] on RoadNet Sequence. Loss weight for n/a tokens are set to zero. Noise cate. stands for noise category.
Thresholds for Reachability Precision-Recall are chosen from $[0.5,1.0,1.5,2.0,2.5]m$.
III-E Auto-Regressive RoadNetTransformer
Based on auto-regressive dependency of RoadNet Sequence, we design our baseline as auto-regressive  RoadNet Sequence generation.
Architecture. We apply the same encoder-decoder architecture as [22]. The encoder is responsible for extracting BEV feature $\mathcal{F}$ from multiple onboard cameras such as Lift-Splat-Shoot [13]. For decoder, we use the same Transformer decoder as [22] which includes a self-attention layer, a cross-attention layer and a MLP layer.
Objective. We denote the ground-truth RoadNet Sequence with length $L$ as $y$ and the predicted RoadNet Sequence as $\hat{y}$, then the objective of auto-regressive RoadNetTransformer is maximum likelihood loss, i.e.
$\max\sum_{i=1}^{L}w_{i}\log{P(\hat{y}_{i}|y_{<i},\mathcal{F})},$ (1)
where $y_{i}$ is the $i^{th}$ token of $y$, $y_{<i}$ means all tokens before $y_{i}$ and $w_{i}$ is the class weight. In practice, since the label Lineal for $v_{c}$ and index 0 for $v_{d}$ appear most frequently, we set $w_{j}$ as a small value for these class.
Input and target sequence construction. The input sequence starts with a start token and the target sequence ends with an EOS token. We also apply synthetic noise objects technique [22] to the sequence construction. Details are shown in the Supplementary material.
Efficiency. Suppose the Transformer spends $\mathcal{T}_{s}$ time inferring a single query. The inference time complexity should be $\mathcal{O}(|\mathcal{E}|\cdot\mathcal{T}_{s})$.
III-F Semi-Autoregressive RoadNetTransformer
The vanilla Auto-Regressive RoadNetTransformer generates RoadNet Sequence one by one, which is a highly time-consuming process. The reason for this expensive one-by-one generation is the ingrained auto-regressive dependency assumption.
In the field of Natural Language Processing, the human language is highly cohesive, which means that any attempt to generate text without auto-regressive assumption can result in a significant decrease in accuracy [50]. However, it’s not the case for road network. With regards to the Figure 1, observations have been made regarding the dependency of RoadNet Sequence: (i) The locations of certain road points (start points, fork points or merge points) can be independent of previous vertices and instead depend solely on the BEV feature map, i.e.,
$P(y_{i}|y_{<i},\mathcal{F})=P(y_{i}|\mathcal{F}).$ (2)
(ii) Except for locations of these road points, other tokens are still strongly auto-regressive.
Drawing from these findings, we suggest the adoption of Semi-autoregressive RoadNetTransformer (SAR-RNTR) that retains auto-regressive functionality within local contexts while simultaneously conducting none-autoregressive generations in parallel. To facilitate this approach, we propose a novel representation of the road graph called Semi-Autoregressive RoadNet Sequence.
Figure 7: Top indicates the selection of key-points and the approach to construct sub-trees taking these key-points as root. Bottom presents the Semi-Autoregressive RoadNet Sequence for the example mentioned above.
Semi-Autoregressive RoadNet Sequence. The objective of Semi-Autoregressive RoadNet Sequence is to divide the trees in the Directed Forest into smaller sub-trees as much as possible so that each tree will be simultaneously inferred and the auto-regressive length can be reduced as much as possible. As shown in Figure 7, we begin by identifying all key-points in the Directed Acyclic Graph (DAG) that meet the condition $od(v)>1$ or $id(v)>1$ or $id(v)=0$. We then proceed to recursively extract these points from their original parent and sub-tree until they become roots with an $id(v)$ value of 0. To restore the edge between the key-point $v$ and its parent, we create a duplicate of the parent and assign it as the child of $v$ with a special label, Clone, identical to that used in RoadNet Sequence. Similarly, the Clone will only be traversed after its original vertex is traversed. As shown in Figure 7, different from the auto-regressive RoadNet Sequence, we construct an independent sequence for each independent tree, so that the SAR-RoadNet Sequence is a 2-dimension sequence, i.e., $y_{1,1},y_{1,2},\cdots,y_{1,L}],\cdots,[y_{M,1},y_{M,2},\cdots,y_{M,L}]]$, where $L$ is the maximum length of each sub-sequence and $M$ is the number of sub-sequences. The padding rules is shown in the Supplementary material. Noted that the new data structure is also a directed forest, therefore the construction and recovering follow all rules in the auto-regressive RoadNet Sequence.
Architecture. The SAR-RNTR can be divided into three parts: (i) Ego-car Feature Encoder, (ii) Key-point Transformer Decoder, (iii) Parallel-Seq Transformer Decoder. Ego-car Feature Encoder follows that in AR-RNTR. Key-point Transformer Decoder is a parallel Transformer decoder [51], which takes a fixed set of learned positional embeddings as input and predict locations of key points based on set prediction [51].
Then, Parallel-Seq Transformer Decoder is proposed for solving mixture of auto-regressive and non-autoregressive problem, i.e..
$\max\sum_{i=1}^{M}\sum_{j=1}^{L}P(y_{i,j}\ |\ y_{1:M,1:j-1},\mathcal{F},\mathcal{V}_{kp}),$ (3)
where $\mathcal{V}_{kp}$ represents location of key-points detected from Key-points Transformer Decoder. For a certain $i$, $y_{i,j}$ is generated auto-regressively, while for a certain $j$, all $y_{1:M,j}$ are generated in parallel. The dependency is illustrates in Figure 4.
However, following this objective will cost $\mathcal{O}(M^{2}\times L^{2})$ memory complexity for self-attention [20]. Inspired by [52], where a cross combination of self-attention from different directions leads to a final global attention, we design two self-attention applied on different directions of the 2-dimension Semi-Autoregressive RoadNet Sequence. As shown in Figure 4, Intra-seq self-attention first applies self-attention on $y_{1:M,j}$ for each $j$ and Inter-seq self-attention then applies self-attention on $y_{i,1:j-1}$ for each $i$. The memory complexity is reduced to $\mathcal{O}(M^{2}+L^{2})\ll\mathcal{O}(M^{2}\times L^{2})$.
Key-point prompt learning. To deduce $y_{i,1:L}$ using $\mathcal{V}_{kp}$ as a basis according to Equation 3, we implement Key-point Prompt Learning. Key-point Prompt for a sub sequence $y_{i,1:L}$ contains two parts: (i) locations of all key-points; (ii) location of the start key-point (Ancestor) of $y_{i,1:L}$. As depicted in Figure 4, this involves organizing the locations of key-points and the start key-point location as a sequence of discrete tokens, and assigning dedicated word embeddings and position embedding for the prompt. The Key-point Prompt is then added to Semi-Autoregressive RoadNet Sequence facilitates the aggregation of key-point information in the sequence.
[…]
The second step is to compute the Hungarian loss for all pairs matched in the previous step. The Hungarian loss is a linear combination of a negative log-likelihood for class prediction and a L-1 loss.
[…]
$+\left.\mathbb{1}_{\{c_{i}\neq 0\}}\|k^{(i)}-\hat{k}^{(i)}\|_{1}\right].$ (4)
Auto-regressive MLE loss follows that of AR-RNTR.
Efficiency and dependency. Suppose the Transformer spends $\mathcal{T}_{s}$ time inferring a single query, and the parallel acceleration rate for GPU is $\alpha\ll 1$. The inference time complexity of combination of Key-point and Parallel-seq Transformer can be approximated as $\mathcal{O}(\alpha(|\mathcal{E}|+|\mathcal{V}_{kp}|)\cdot\mathcal{T}_{s})$ which greatly less than that of Auto-Regressive by ratio $\alpha(1+\frac{|\mathcal{V}_{kp}|}{|\mathcal{E}|})$. In addition to acceleration, improved dependency modeling can also enhance contextual reasoning [53, 54, 55], thus benefiting performance.
Figure 8: Left illustrates the training stage which takes masked (gray square) Semi-Autoregressive RoadNet Sequence as input. Right shows inference stage where the Parallel-seq Transformer Decoder takes a fully-masked sequence as input and iteratively masks the predicted token with low confidence.
III-G Non-Autoregressive RoadNetTransformer
While the Semi-autoregressive RoadNetTransformer improves inference speed to some extent, there is still an inefficient auto-regressive component present. Therefore, we propose a fully non-autoregressive generation model, called Non-Autoregressive RoadNetTransformer, which can output the entire sequence at once. To mimic the auto-regressive generation, we iteratively refine the output from this non-autoregressive generation , i.e.,
$\max\sum_{i=1}^{M}\sum_{j=1}^{L}P(y_{i,j}\ |\ \hat{y},\mathcal{F},\mathcal{V}_{kp}),$ (5)
where $\hat{y}$ is the predicted Semi-Autoregressive RoadNet Sequence of the last guess. Provided the limited iteration times, the Non-Autoregressive RoadNetTransformer can achieve a extremly high inference speed than its auto-regressive counterpart.
The training and inference approach are visualized in Figure 8. During training, we utilize a masked language modeling strategy [56] that involves masking a high percentage of the input ground-truth sequence and prompting the Transformer to predict all the missing tokens. Thus, during inference, we begin with a fully masked sequence and predict the Semi-Autoregressive RoadNet Sequence multiple times, with each iteration masking tokens with low confidence. With each iteration, the results will be gradually refined.
Efficiency. Time complexity for NAR-RNTR can be approximated as $\mathcal{O}(\alpha|\mathcal{V}_{kp}|(N_{iter}+1)\cdot\mathcal{T}_{s})$. Acceleration ratio is $\frac{|\mathcal{E}|/|\mathcal{V}_{kp}|+1}{N_{iter}+1}$ where $N_{iter}\ll|\mathcal{E}|/|\mathcal{V}_{kp}|$.
III-H Details of RoadNet Sequence Representation
Our method first formalizes road networks as directed acyclic graphs (DAGs) encoding both Euclidean geometry and non-Euclidean topology (Section III (A)). These graphs are then serialized into two types of representations: (i) a coupled format, which binds geometry and connectivity tightly (Section III (B)), and (ii) a decoupled format, which separates node and edge information for modeling flexibility (Section III (C)). To decode these sequences, we design three Transformer-based strategies: (i) an autoregressive decoder (AR-RNTR) that generates tokens sequentially (Section III (E)), (ii) a semi-autoregressive decoder (SAR-RNTR) that enables partial parallel decoding through keypoint grouping (Section III (F)), and (iii) a non-autoregressive decoder (NAR-RNTR) that predicts all tokens in parallel and refines them iteratively (Section III (G)). For the decoding process, we represent each vertex–edge pair using six integers: the vertex location ($v_{x}$, $v_{y}$), vertex category $v_{c}$, index of the parent vertex $v_{d}$, and the Bezier curve coefficients ($e_{px}$, $e_{py}$). The details are provided in Table I. Together, as illustrated in Figure 9, these components collectively contribute to a cohesive and robust representation that significantly enhances road network construction and decoding process in terms of geometry, topology, and computational efficiency.
Recognizing that perception errors from BEV inputs can propagate and mislead topology generation, we introduce two enhancement modules. The first is a topology-inherited training (TIT) strategy that decouples perception and reasoning by pretraining topology on clean BEV maps, distilling it to noisy inputs, and fine-tuning jointly, as shown in Section III (K). The second enhancement leverages SD-Map sequences as prompts, injecting prior knowledge of road connectivity to guide generation and boost generalization, as shown in Section III (L). Together, as in Figure 10, these components form a cohesive and robust system that significantly improves road network reconstruction across geometry, topology, and efficiency dimensions.
Figure 9: A visual explanation of RoadNet sequence construction and decoding process.
Figure 10: A visual illustration of the entire method, including the proposed RoadNet sequence, the three models, and the enhancement modules used to improve performance.
Figure 11: Examples of TP, FP and FN when evaluating reachability between $v_{1}$ and $v_{4}$. Although case 1 misses landmark $v_{2}$, but both paths between $v_{1}$ and $v_{4}$ are within distance of the ground-truth paths. Case 2, however, has path 2 surpasses the threshold distance from matched its matched ground-truth path 2 so it’s false positive. Case 3 predicts path 1 twice but they are both within threshold distance. However, path 2 in the ground-truth has no matched prediction so it’s false negative.
III-I Metrics
Precision-recall. Precision is defined as
$\text{Precision}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Positive}}.$ (6)
Recall is defined as
$\text{Recall}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Negative}}.$ (7)
And F1 score is defined as
$\text{F1 score}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}.$ (8)
To ensure a fair comparison, we employ three metrics from [17, 18]: mean precision-recall, detection ratio, and connectivity. However, these metrics, which rely solely on centerline detection, neglect the significance of both the location accuracy of road-points and the reachability of the road graph. To make up with the deflect, we propose 2 following metrics.
Landmark precision-recall. We use landmark precision-recall to evaluate the location accuracy of landmarks. For each predicted landmark, we match it to a ground-truth with the nearest distance. If a predicted landmark is within the threshold distance with its matched ground-truth, it is true positive, otherwise it is false positive. If a ground truth landmark is not matched with any predictions or not within the threshold distance with its matched prediction, it is false negative. Thresholds for landmark precision-recall are chosen from $[0.5,1.0,1.5,2.0,2.5,$ $3.0,3.5,4.0,4.5,5.0]m$.
Reachability precision-recall. One of the motivations for predicting road topology is finding the valid paths of any two points on the road, which allows the self-driving car to reach its destination. Reachability between landmarks is defined that there exists a path going from one landmark to another. After matching landmarks in the former session, we propose reachability precision-recall to evaluate both connectivity and accuracy of path between landmarks. In the road network, if two landmarks are connected, it means there exists a path from one landmark to another. Given a pair of predicted landmarks $\hat{A},\hat{B}$ and its matched pair of ground-truth landmarks $A,B$, a true positive is a path connecting $\hat{A},\hat{B}$ with Chamfer Distance to any of the ground-truth path connecting $A,B$ less than the threshold. A false negative is a ground truth path from $A,B$ with no matched predicted path. Thresholds for reachability precision-recall are chosen from $[0.5,1.0,1.5,2.0,2.5]m$. An example is given in Figure 11.
Figure 12: Visualization of road segmentation and road network extraction results on the new split of the nuScenes [24] dataset. Top: The misprediction of a crossroad obstructed by a barrier impairs the AR-RNTR’s ability to predict its topology (circled in orange). Bottom: The misprediction of a U-turn and T-road obstructed by terrain also harms the topology prediction.
Figure 13: Left: In the first step, a simple ResNet [57] is used as the BEV Encoder to encode the BEV Lidar map into BEV features, and RoadNetTransformer is trained using the clear BEV features from the Lidar map. Middle: In the second step, the BEV Encoder with the Lidar map input is frozen, and we distill the BEV Encoder using round-view images as input to match the previously obtained BEV features. Right: Finally, the BEV Encoder using round-view images as input and RoadNetTransformer are assembled and fine-tuned for a few epochs.
III-J Bottleneck of Sequence-to-sequence Methods
After proposing our auto-regressive, semi-autoregressive, and non-autoregressive sequence-to-sequence methods for road network extraction, we examine the bottlenecks of the sequence-to-sequence approach. First, since the default split of the nuScenes dataset includes overlapping scenes in the training and evaluation sets, we need to create a new split to ensure non-overfitting. Unlike before [25], we should also ensure that different daytime and weather conditions are equally distributed in the training and evaluation sets. We then retrain the road segmentation model on the new split dataset and train the AR-RNTR. The visualization of the results is presented in Figure 12. Clearly, occlusion in the round-view camera cannot be avoided, making accurate road perception impossible even in ideal conditions. For example, the far side of a crossroad or a U-turn cannot be predicted as they are invisible. Therefore, in the AR-RNTR results, the inaccurate Euclidean understanding impairs the perception of non-Euclidean data, meaning location errors propagate to the topology learning.
Since the Euclidean understanding is primarily the responsibility of the BEV encoder, decoupling the training of the BEV encoder and RoadNetTransformer is a natural choice to mitigate error propagation during training. First, RoadNetTransformer is trained to enhance its topological understanding, and then we transfer the topology knowledge to both the BEV encoder and RoadNetTransformer. We name this new training strategy Topology-Inherited Training (TIT).
III-K Topology-Inherited Training
As mentioned earlier, the Euclidean understanding is primarily the responsibility of the BEV Encoder, so we need to decouple the training of the BEV Encoder and RoadNetTransformer. However, the training of RoadNetTransformer also depends on the BEV features as input. Therefore, we utilize the Lidar map provided by the nuScenes [24] dataset, which contains high-resolution Birds-Eye View (BEV) road images as an intermediate step. As shown in the first column of Figure 13, we first use a simple ResNet [57] as the BEV encoder for the BEV Lidar map, and RoadNetTransformer uses the clear BEV features as input to train the RoadNet Sequence.
Since the first stage is crucial and much more efficient than using round-view images as input, we utilize the entire nuScenes [24] dataset with a 20Hz sample rate instead of the 2Hz rate. This approach increases the scale of data to a factor of nine.
TIT decomposes the original objective function
$\min_{\theta_{f},\theta_{g}}\mathbb{E}_{(x,y)\sim D}\mathcal{L}_{\text{CE}}(g(f_{I}(x)),y),$ (9)
into a two-step optimization process:
$\min_{\theta_{g}}\mathbb{E}_{(x_{L},y)\sim D_{L}}\mathcal{L}_{\text{CE}}(g(f_{L}(x_{L})),y),$ (10)
$\min_{\theta_{f}}\mathbb{E}_{(x,x_{L})\sim D}\|f_{I}(x)-f_{L}(x_{L})\|_{1}.$ (11)
This progressive optimization strategy mitigates gradient oscillations caused by directly optimizing a complex objective, thereby enhancing stability and accelerating overall convergence.
In the second stage, as shown in the second column of Figure 13, we train the BEV Encoder with round-view image input using distillation. We freeze the parameters of the BEV Encoder when using the Lidar map input and keep the parameters open for training when using the round-view images input. We employ L-1 loss, so the loss function $\mathcal{L}_{dist}$ can be expressed as:
$\mathcal{L}_{dist}=\|\mathcal{F}_{img}-\mathcal{F}_{lidar}\|_{1}.$ (12)
Compared to directly optimizing the BEV encoder using cross-entropy loss for the final objective, L1 loss provides more stable gradients, mitigating gradient oscillations caused by topology prediction errors in end-to-end training. This gradient smoothness ensures more stable parameter updates for the BEV encoder, reducing training instability caused by large gradient fluctuations.
In the final stage, we assemble the BEV Encoder with round-view image input and RoadNetTransformer as in the traditional AR-RNTR method, and train for a few epochs to fine-tune the model.
Analysis. To minimize error propagation in topology reasoning, we adopt a three-stage approach. In the first stage, we use Lidar data to generate accurate BEV features, which are then used to train the RoadNet Transformer for RoadNet Sequence learning. In the second stage, we distill LiDAR-based BEV features to camera-based BEV features to refine the encoder’s Euclidean understanding. Finally, in the third stage, we integrate the camera-based BEV features with the RoadNet Transformer for joint learning. This decoupling allows each component to focus on its specific task—accurate BEV feature generation and RoadNet Sequence learning—effectively reducing error propagation throughout the training process.
The BEV Encoder faces several challenges that impact its effectiveness in complex driving scenarios. First, due to its fixed-resolution grid representation, it may miss fine details of road networks, particularly in dense urban areas where lane markings and road boundaries are closely packed. The discretization process results in information loss, which can degrade spatial precision and reduce the encoder’s ability to capture fine-grained structural and semantic features necessary for accurate perception and planning. Second, the encoder struggles with non-uniform road layouts, especially in multi-level structures like overpasses and tunnels. Since BEV projections compress 3D spatial information into a 2D plane, it becomes difficult to distinguish overlapping roads, leading to potential connectivity errors in topology reasoning. Third, the reliance on 2D projections limits the encoder’s ability to interpret depth and elevation, making it less effective on uneven terrains. To mitigate these issues, enhancements such as higher-resolution BEV features, multi-scale feature aggregation, and 3D-aware representations could improve accuracy and generalization.
Figure 14: Illustration of the SD-Map data collection pipeline. First, we select the area on OpenStreetMap [27] and export its topology data. Then, we manually align the SD-Map with the nuScenes [24] map. Since the two maps have some local distortions, we manually adjust some nodes to achieve approximate local alignment.
Figure 15: Comparison between SD-Map and road network in several examples.
III-L Standard Definition Map (SD-Map)
Although TIT partly solves the bottlenecks, the obstacles of roads from barriers or terrain shown in Figure 12 are not fundamentally resolved. Therefore, to address this problem, some prior information should be provided. The prior information should be easy to obtain and inexpensive to process.
In daily driving experiences, navigation maps like Google Maps [26] are almost a necessity worldwide, making it easy to gather some information from them. As shown in Figure 2, information in these navigation maps is abundant and difficult for algorithms to decode. Therefore, we extract only road topology information from the navigation map, commonly referred to as standard definition maps (SD-Maps), as shown in the middle of Figure 2.
The comparison between SD-Map and road network is shown in Figure 15. Unlike road network, which provides lane-level and centimeter-level annotations, SD-Maps contain only road-level and meter-level information. This means they cannot provide the accurate number of lanes for each road, and the locations of landmarks are not precise. Additionally, the road-level annotations imply that some edges between vertices are bi-directional, which means the road network cannot be abstracted as a Directed Acyclic Graph (DAG) since cycles may exist.
Collection of SD-Map. We collect SD-Map data from OpenStreetMap [27], which is a free and open geographic database based on freely licensed geodata sources. We export Tracktopo data from the database, whose raw data appears as shown in the second step of Figure 14. The raw SD-Map is constructed from nodes and their connections with direction, which is very similar to road network. Next, we align the SD-Map with the nuScenes map. Since the nuScenes map does not provide GPS coordinates, we must align them manually. We also manually adjust some nodes to correct obvious errors in the alignment. It should be noted that this human intervention is due to the deficiencies of the nuScenes dataset. In practice, acquiring SD-Map information from navigation maps is easy and efficient. Since we do not expect SD-Maps to provide highly accurate locations, some location errors are acceptable.
Construction of SD-Map Sequence. Since the SD-Map is also constructed with vertices and edges, we can similarly project SD-Maps to sequences as we have done with RoadNet Sequence. The only difference is that SD-Maps may not be Directed Acyclic Graphs since cycles may exist. Therefore, we first need to transform the cyclic graph into a DAG. To achieve this, we duplicate the vertex whenever it is traversed again during a Depth-First Search. The new vertex will copy the location of its original vertex and be noted as the leaf vertex of the last traversed vertex. After being transformed into a DAG, we can simply use coupled or decoupled RoadNet Sequence to obtain what we call SD-Map Sequence. To be noted, to reduce the length of SD-Map Sequence, we do not include the Bezier curve control points and only treat edges as straight lines.
SDMap-RoadNetTransformer. Based on the success of Auto-regressive RoadNetTransformer (AR-RNTR), which treats road network extraction as a sequence-to-sequence approach, we can simply upgrade to SDMap-RoadNetTransformer (SD-RNTR). We use SD-Map Sequence as a sequence prompt before the start of RoadNet Sequence. This change does not alter the AR-RNTR model. Without modifying the structure of AR-RNTR or undergoing complex data processing for a new kind of data, RoadNetTransformer can easily utilize the SD-Maps, resulting in a significant performance boost.
IV Experiments
IV-A Dataset
We utilize the nuScenes [24] dataset (700/150/150 for training/validation/test) to assess our approach. Regardless of the split type, the number of samples in each set remains unchanged. We only utilize sensor information from six cameras, IMU, and GPS, which were sampled at a rate of 2Hz. When training RoadNetTransformer in the first stage of Topology-Inherited Training, we utilize sensor information from IMU and GPS sampled at a rate of 20Hz.
To demonstrate the generalization ability of our approach, we further evaluate it on the Argoverse [58] dataset. The Argoverse dataset is a large-scale autonomous driving benchmark comprising diverse multimodal data collected from real-world urban driving in Pittsburgh and Miami. It contains 113 sequences with over 290,000 annotated 3D object tracks, covering approximately 5 million object instances. The dataset is complemented by high-definition maps spanning nearly 290 kilometers of road networks, which include lane geometry, drivable areas, and traffic semantics, enabling rich contextual modeling.
IV-B Implementation
Pretrain. We follow LSS [13] for BEV Encoder. The input images are resized to $128\times 352$, and the target BEV area is from -48 to 48m in x-direction (roll) and -32 to 32m in y-direction (pitch) with resolution 11m in the ego coordinate system. ResNet-50 [57] or VoVNetV2 [59] are used as the image backbone. The ResNet-50 [57] model is pretrained using DeepLabV3+ [60] on the Cityscapes dataset[61], and the VoVNetV2 is pretrained on FCOS3D [62]. To be noted that the VoVNetV2 is pretrained on the default split of the nuScenes [24] dataset; therefore, we will not use the VoVNetV2 in the non-overfitting split. Initially, we pretrain the BEV encoder on centerline segmentation, following all the training strategies outlined in LSS [13]. During the training of RoadNetTransformer, we load and freeze the parameters of the BEV encoder.
RoadNetTransformer. Details of three variants are shown below. AR-RNTR: The length of RoadNet Sequence is padded to $6\times 100$. Transformer decoder layers of AR-RNTR are set as 6. Our AR-RNTR is trained on 300 epochs with learning rate $2\times 10^{-4}$, batch size as $2\times 8$. During the training process, we apply random flip, random rotation, and random scaling on BEV feature similar to [29, 16]. SAR-RNTR: Transformer decoder layers of SAR-RNTR are set as 6 (Key-point) plus 3 (Parallel-seq). We use noise vertices to pad all sub-sequences to the length of $6\times 18$, and the max number of key-points is set to 34. The strategies for data augmentation and training are the same as AR-RNTR. NAR-RNTR: During training, we load parameters of trained SAR-RNTR and then finetune it with a mask language modeling strategy for 100 extra epochs. During inference, the iteration number is set to 3.
Topology-Inherited Training. We apply Topology-Inherited Training only to AR-RNTR. The basic settings of AR-RNTR are the same as those previously introduced for RoadNetTransformer. In the first stage, we train on 20Hz data for 300 epochs. We use more data augmentation than when training RoadNetTransformer. The random rotation is increased from $45^{\circ}$ to $45^{\circ}$, and the scale is adjusted from 0.9 to 1.1. In the second stage, we train for 24 epochs. In the third stage, we train for 100 epochs.
SD-Map. We locate Singapore onenorth, Singapore Queenstown, Singapore Hollandvillage, and Boston Seaport from OpenStreetMap [27] and approximately choose the corresponding areas. After exporting the topology data, we retain only the nodes of highway and their secondary type of link roads.
Metrics. Thresholds for Mean Precision-Recall and Landmark Precision-Recall are uniformly sampled from $[0.5,5.0]$ m. Thresholds for Reachability Precision-Recall are uniformly sampled from $[0.5,2.5]$ m. We only take into account the reachability between vertices that are within a maximum of 5 edges of connection.
IV-C Results
TABLE II: Comparison of front-camera road network extraction with state of the art on nuScenes [24] PON validation split [25]. ResNet-50 [57] is applied as image backbone by default. M-P, M-R, M-F stand for mean precision/recall/F1-score [17]. Detect stands for Detection ratio metrics [17]. C-P, C-R, C-F stand for connectivity precision/recall/F1-score [17].
Methods M-P M-R M-F Detect C-P C-R C-F
PINET [63] 54.1 45.6 49.5 19.2 - - -
Poly [19] 54.7 51.2 52.9 40.5 58.4 16.3 25.5
STSU [17] 60.7 54.7 57.5 60.6 60.5 52.2 56.0
TPLR [18] - - 58.2 60.2 - - 55.3
AR-RNTR 60.9 57.9 59.3 61.7 63.2 52.7 57.5
SAR-RNTR 63.5 59.9 61.6 63.5 67.1 57.2 61.7
NAR-RNTR 62.0 59.4 60.7 62.0 66.4 56.2 60.9
TABLE III: Comparison on Argoverse [58] dataset. ResNet-50 [57] is applied as image backbone by default. M-F and C-F stand for mean and connectivity F1-score [17]. Detect stands for detection ratio metrics [17].
Methods M-F Detect C-F
PINET [63] 47.2 15.1 -
TR [17] 55.6 60.1 54.9
TPLR [18] 57.1 64.2 58.1
Ours (AR-RNTR) 57.6 64.5 58.2
Comparison with state of the art. We compare our model with previous state-of-the-art methods on high-definition road network topology extraction. To achieve a fair comparison, we only utilize front-view images as input and ResNet-50 [57] trained on ImageNet-1K [64] as the backbone. Also, we utilize the PON nuScenes train/val split [25] as also applied in [17, 18]. Table II presents the results of our approach on the nuScenes dataset using the Mean Precision-Recall, Detection, and Connectivity metrics [17]. On the one hand, our model outperforms all three variants across all metrics, demonstrating the superior performance of RNTR in both centerline detection and centerline association estimation. This remarkable improvement on both centerline location and connectivity can be attributed to the unified representation provided by RoadNet Sequence and the global context reasoning capabilities of the Transformer architecture. We also validate our method on Argoverse dataset in Table III. The experimental results demonstrate that our method consistently outperforms the competitors.
On the other hand, unlike in Natural Language Processing, our Semi-Autoregressive and Non-Autoregressive versions significantly outperform the Auto-Regressive version, highlighting our dependency decoupling. sc
TABLE IV: Comparison of multiple-camera road network extraction on nuScenes [24] dataset assessed by Landmark Precision-Recall (L-P, L-R, L-F) and Reachability Precision-Recall (R-P, R-R, R-F). ResNet-50 [57] is applied as image backbone by default. $\star$” use coupled road network. “†” use VoVNetV2 [59]
Methods Landmark Reachability FPS
L-P L-R L-F R-P R-R R-F
[…]
Split Methods Landmark Reachability
L-P L-R L-F R-P R-R R-F
[…]
AR-RNTR 62.5 47.7 54.1 73.2 52.8 61.3
SAR-RNTR 65.3 49.1 56.0 74.1 59.1 65.7
NAR-RNTR 64.1 48.3 55.1 74.0 58.7 65.1
AR-RNTR + TIT 67.7 53.6 59.8 79.3 63.8 70.7
SD-RNTR + TIT 77.1 69.4 73.0 75.9 68.9 72.2
New AR-RNTR 47.7 29.6 36.5 52.1 33.6 40.9
SAR-RNTR 50.3 29.2 36.9 55.9 35.0 43.0
NAR-RNTR 48.7 28.8 36.2 55.6 34.1 42.3
AR-RNTR + TIT 48.8 32.7 39.1 67.1 39.6 49.8
SD-RNTR + TIT 63.0 48.3 54.7 67.9 58.1 62.6
Over-fitting. Table V shows results with the new split on the nuScenes dataset. Compared to the default split, the results from the non-overlapping split show a significant reduction in both Landmark and Reachability metrics. The decrease in Landmark performance aligns with the observation that without overfitting, invisible roads cannot be predicted through round-view images. Comparing the three variants of RoadNetTransformer, the Landmark scores show no significant differences, all hovering around 36.5, suggesting that landmark detection has reached its upper limit. In non-overlapping scenes, the primary challenge is the detection of roads, which leads to less accurate landmark detection compared to the overfitting scenario. This inaccuracy also adversely affects Reachability.
Topology-Inherited Training (TIT.) After applying Topology-Inherited Training, AR-RNTR experiences a significant improvement in both the overfitting and non-overfitting splits. In the non-overfitting split, TIT helps AR-RNTR exceed the performance of both SAR-RNTR and NAR-RNTR. Especially in Reachability, TIT leads SAR-RNTR by 11.6, demonstrating that with the help of topology data inheritance, topology training is more effective. In terms of Landmark, however, the improvement is not as obvious, with less than a 3-point increase from AR-RNTR. This indicates that TIT cannot fundamentally improve road landmark detection.
SD-Maps. The usage of SD-Maps as prior information brings significant improvements for both the overfitting and non-overfitting splits, with more pronounced benefits for the non-overfitting split. It raises the Landmark F-score from 39.1 to 54.7 and the Reachability F-score from 49.8 to 62.6. The most notable improvement in the Landmark F-score demonstrates the effectiveness of SD-Maps. Additionally, better landmark detection supports improved topology detection.
IV-D Ablation Studies
We conduct ablation studies on Transformer layer numbers, as well as mask ratio and iteration times of NAR-RNTR during inference.
TABLE VI: Ablation studies on number of Transformer decoder layers and Intra-seq self-attention of NAR-RNTR. # Key-pt denotes Key-point Transformer Decoder layer number, and # Para-seq denotes Parallel-seq Transformer Decoder layer number. Intra-Seq means Intra-seq self-attention in Parallel-seq Transformer Decoder layer. VoVNetV2 [59] pretrained on extra data is applied as image backbone by default. The row with gray color is our final choice.
# Key-pt # Para-Seq Intra-Seq L-F R-F FPS
3 3 56.1 63.3 4.8
5 3 58.2 65.1 4.8
6 3 60.2 66.0 4.7
6 5 60.5 66.9 3.4
6 6 61.3 67.1 3.0
6 3 55.7 56.3 5.0
Number of Transformer layers. We investigated the impact of the number of Transformer layers on NAR-RNTR in Table VI. Using fewer Key-point Transformer decoder layers leads to a significant loss in accuracy for both landmark and reachability, but provides a limited speedup. Conversely, using fewer Parallel-seq Transformer decoder layers dramatically improves inference speed while incurring less accuracy loss.
Intra-Seq self-attention. The bottom row of Table VI highlights the crucial role of Intra-seq self-attention. Its absence results in a significant loss in both landmark localization and topology connection.
TABLE VII: Ablation studies on mask ratio and iteration times of NAR-RNTR. VoVNetV2 [59] pretrained on extra data is applied as image backbone by default. The row with gray color is our final choice.
Mask ratio # iteration L-F R-F FPS
50% 3 51.4 51.3 4.7
75% 3 59.4 62.7 4.7
90% 1 59.0 62.7 8.9
90% 3 60.2 66.0 4.7
90% 6 60.3 66.1 2.8
Mask ratio and iteration times. Our investigation into the impact of the mask ratio during the training of NAR-RNTR revealed the importance of using a large mask ratio. Consequently, a 90% mask ratio is used during Non-Autoregressive finetuning. We also observed accuracy saturation as the number of iterations increased. Therefore, we used 3 iterations for our approach.
TABLE VIII: Ablation studies on BEV augmentation and synthetic noise objects [22] (including sequence augmentation and sequence noise padding). NAR-RNTR with VoVNetV2 [59] pretrained on extra data is applied. The row with gray color is our final choice.
BEV Aug Sequence Aug Sequence Noise L-F R-F
58.6 64.3
57.5 62.7
60.2 66.0
59.1 65.2
TABLE IX: Ablation studies on sequence embedding size and class weight. NAR-RNTR with VoVNetV2 [59] pretrained on extra data is applied. The row with gray color is our final choice.
Embedding size class weight L-F R-F
576 1.0 60.1 65.5
576 0.5 60.1 66.1
576 0.1 60.2 66.0
576 0.2 60.2 66.0
1000 0.2 60.1 65.8
2000 0.2 59.7 65.5
TABLE X: Ablation studies on Topology-Inherited Training. Models with ResNet [57] trained on the nuScenes [24] non-overfitting dataset are used. The ”Lidar Map” stage indicates the first stage of TIT, and ”Assembled” represents the third stage of TIT. ”Sweep data” uses 20Hz sampling rate data from the nuScenes dataset.
Stage Methods Landmark Reachability
L-P L-R L-F R-P R-R R-F
Lidar Map Baseline 47.8 38.9 42.9 60.4 24.9 35.3
+ More aug 52.6 41.4 46.3 70.9 25.6 37.6
+ Sweep data 67.2 47.2 55.4 81.1 51.0 62.6
Assembled Freeze BEV 47.7 34.2 38.6 66.7 38.2 48.6
Unfreeze BEV 48.8 32.7 39.1 67.1 39.6 49.8
TABLE XI: Comparison on nuScenes [24] dataset assessed by Landmark Precision-Recall (L-P, L-R, L-F) and Reachability Precision-Recall (R-P, R-R, R-F). The model used is AR-RNTR. ResNet-50 [57] is applied as image backbone by default.
Learning Rate Landmark Reachability
L-P L-R L-F R-P R-R R-F
1e-4 61.0 46.8 53.0 72.5 52.1 60.6
2e-4 62.5 47.7 54.1 73.2 52.8 61.3
3e-4 62.0 47.2 54.0 72.8 52.4 60.7
TABLE XII: Comparison on nuScenes [24] dataset assessed by Landmark Precision-Recall (L-P, L-R, L-F) and Reachability Precision-Recall (R-P, R-R, R-F). ResNet-50 [57] is applied as image backbone by default.
Methods Landmark Reachability
L-P L-R L-F R-P R-R R-F
[…]
TABLE XIII: Comparison of different geographical location settings on nuScenes [24] PON validation split [25]. ResNet-50 [57] is applied as image backbone by default. M-P, M-R, M-F stand for mean precision/recall/F1-score [17]. Detect stands for Detection ratio metrics [17]. C-P, C-R, C-F stand for connectivity precision/recall/F1-score [17].
Method nuScenes
Singapore Boston
M-P M-R M-F Detect C-P C-R C-F M-P M-R M-F Detect C-P C-R C-F
AR-RNTR 61.1 57.8 59.6 61.8 63.1 52.5 57.9 60.8 58.2 59.1 61.5 63.4 52.8 57.3
SAR-RNTR 63.8 60.4 62.0 63.4 67.3 56.8 61.2 63.0 59.6 61.1 63.5 66.8 57.5 62.0
NAR-RNTR 61.4 59.1 60.2 62.3 66.8 55.7 60.5 62.5 59.6 61.1 61.8 66.2 56.4 61.2
TABLE XIV: Comparison of different geographical location settings on Argoverse [58] dataset. ResNet-50 [57] is applied as image backbone by default. M-F and C-F stand for mean and connectivity F1-score [17]. Detect stands for detection ratio metrics [17].
Method Argoverse
Miami Pittsburgh
M-F Detect C-F M-F Detect C-F
AR-RNTR 58.2 64.8 58.4 57.2 64.1 58.1
SAR-RNTR 60.1 65.4 61.2 60.6 66.3 62.0
NAR-RNTR 58.4 65.6 60.7 57.6 64.3 60.2
TABLE XV: Comparison of different weather condition settings on nuScenes [24] PON validation split [25]. ResNet-50 [57] is applied as image backbone by default. M-F stand for mean F1-score [17]. Detect stands for Detection ratio metrics [17]. C-F stand for connectivity F1-score [17].
Method nuScenes
Normal Rainy Cloudy
M-F Detect C-F M-F Detect C-F M-F Detect C-F
AR-RNTR 59.9 62.8 58.4 58.6 61.3 56.9 59.0 61.5 57.3
SAR-RNTR 62.4 63.9 62.6 60.7 63.1 61.1 61.8 63.6 62.0
NAR-RNTR 62.0 63.5 62.1 60.3 61.5 60.3 60.2 61.7 60.4
Non-unique sorting. We show the difference between the random ordering strategy and an ordering based on coordinates in Table
BEV augmentation. The first column of Table VIII shows that the BEV augmentation provides a significant 2.7/3.3 improvement in both Landmark and Reachability scores.
Synthetic noise objects. The second column of Table VIII shows that the sequence augmentation of Synthetic noise objects technique [22], however, leads to a drop in performance. Whereas, the third column shows that the sequence noise padding 1.6/1.7 improved on both Landmark and Reachability scores. But the sequence noise padding is less effective than BEV augmentation.
Class weight. We exam the class weight for MLE loss, i.e., $w$ for
$\max\sum_{i=1}^{L}w_{i}\log{P(\hat{y}_{i}|y_{<i},\mathcal{F})}.$ (13)
$\max\sum_{i=1}^{M}\sum_{j=1}^{L}w_{j}P(y_{i,j}\ |\ \hat{y},\mathcal{F},\mathcal{V}_{kp}).$ (14)
Due to the high frequency of Lineal for $v_{c}$ and the default index for $v_{d}$, we assign a lower weight to these categories. Although the second column of Table IX does not indicate a clear relationship between class weight and performance, using a lower weight for the loss results in more stable performance.
Embedding size. If we extend the embedding size from 576 to 1000 or 2000, useless embeddings clearly harm the performance.
Topology-Inherited Training. In Table X, we present the ablation study for Topology-Inherited Training (TIT). In the first stage, we train using the lidar map as input. For the baseline, we use the same settings as AR-RNTR, but change the input to the lidar map. Since the input is simpler this time, adding more intense augmentation can significantly reduce overfitting. However, compared to AR-RNTR in Table V, the Landmark score shows a noticeable improvement due to the easier input modality. Nonetheless, the reachability performance is not as good as with the round-view image. The potential problem is overfitting with this simple modality, so we use the sweep data, which is ten times the sample dataset. After training on the sweep data, both the landmark and reachability metrics achieve the best results.
After we assemble LSS and RoadNetTransformer in the third stage of TIT, we study the difference between freezing and unfreezing the BEV Encoder. The difference is not significant, indicating that the distillation in stage 2 already aligns well with the training of RoadNetTransformer. However, fine-tuning the BEV Encoder can slightly better match these two components.
Parameter fine-tuning. We conducted experiments on hyperparameter tuning by varying the learning rate. The results, presented in Table XI, show that our model performs best with the default setting and remains stable with minimal variation when the learning rate is adjusted.
Computational costs. Besides the FPS reported in Table IV, we further provide details on the model parameters and inference time. The base model, AR-RNTR, has 37.25 million parameters, while NAR-RNTR has 35.90 million parameters and requires approximately 6GB of memory during inference with a batch size of 2, where the input images are 128 × 352.
Standard deviations. Including standard deviations strengthens the reliability of our findings. We have provided the mean and standard deviation for AR-RNTR, SAR-RNTR, and NAR-RNTR, along with the number of trials. Each experiment was repeated five times, and the results are presented in Table XII. The performance of our models is statistically stable, with minimal variation.
Figure 16: Mean Precision/Recall v.s. thresholds. $\dagger$}” use VoVNetV2 [59] pretrained on extra data as backbone. Thresholds are from [0.5,1.0,1.5,2.0,2.5,3.0,3.5,4.0,4.5,5.0]m[0.5,1.0,1.5,2.0,2.5,3.0,3.5,4.0,4.5,5.0]m.
IV-E Precision-Recall Curve
In addition to our overall advantage in mean Precision-Recall (as presented in Table 1 of the main submission), Figure 16 displays the precision/recall versus thresholds curve. Our models outperform others in terms of precision and recall for smaller thresholds, highlighting our accuracy advantage.
IV-F Robustness
To evaluate the robustness of our method, we group the evaluation results by geographic region and report performance separately. As shown in Table XIII and Table XIV, our model consistently performs well across diverse urban environments. Specifically, in the nuScenes dataset, we observe strong and stable performance in both Singapore and Boston, demonstrating generalization across cities with different road structures and traffic styles. Similarly, on the Argoverse dataset, which contains data from Miami and Pittsburgh, our method also shows solid performance, indicating good adaptability to diverse North American cities.
Furthermore, to assess robustness under varying weather conditions, we conduct additional experiments on the nuScenes validation split by categorizing scenes based on annotated weather labels (i.e., normal, rainy, and cloudy). As reported in Table XV, our method maintains strong performance across all weather conditions. This confirms the robustness and reliability of our model under challenging real-world scenarios.
Overall, the results across these three tables clearly demonstrate that our approach is resilient to variations in both geography and weather, supporting its effectiveness in real-world deployment.
V Conclusion and Discussions
Conclusion. In summary, our work introduces a lossless, efficient, and interactional sequence representation called RoadNet Sequence, which preserves both Euclidean and non-Euclidean data of road networks. We have designed an Auto-Regressive RoadNetTransformer as a baseline model, which takes advantage of the auto-regressive dependency of RoadNet Sequence. Additionally, we have proposed Semi-Autoregressive and Non-Autoregressive RoadNetTransformer models, which decouple the auto-regressive dependency of RoadNet Sequence, resulting in significantly faster inference speeds and improved accuracy. Our extensive experiments demonstrate the superiority of our RoadNet Sequence representation and RoadNetTransformer models.
To further adapt RoadNetTransformer for real-world applications, we conduct more experiments on the non-overfitting split of the nuScenes [24] dataset. We identify two main bottlenecks in the current RoadNetTransformer: poor landmark detection limited by the BEV Encoder and error propagation to topology reasoning. Therefore, we propose Topology-Inherited Training to inherit better topology knowledge into RoadNetTransformer. Additionally, we collect SD-Maps from open-source map datasets and use this prior information to significantly improve landmark detection and reachability.
Limitations. While our method demonstrates strong performance across various metrics and datasets, there are still some limitations that deserve future exploration. First, our current BEV Encoder relies on 2D projections, which may struggle to capture elevation or fine-grained vertical structures (e.g., tunnels, flyovers), especially in complex 3D scenes. Second, while SD-Maps offer valuable priors, their coarse resolution limits their effectiveness for precise landmark localization, which may restrict performance in high-precision navigation scenarios. We believe these challenges provide promising directions for future work, including exploring 3D aware representations, and leveraging richer map priors with higher spatial fidelity.